OpenStudy (anonymous):

HELLPPP!! what's an explicit definition for the sequence -3, 1, 5, 9 ?!

OpenStudy (anonymous):

what do you mean by explicit definition?

OpenStudy (anonymous):

A formula that allows direct computation of any term for a sequence a1, a2, a3, . . . , an, . . . .

OpenStudy (anonymous):

nth term= a+ (n-1)d where a= -3 d= 4

OpenStudy (anonymous):

for ex 10th term= -3 + (9)4 = 33 Is that what you are looking for?

OpenStudy (the_fizicx99):

You want an explicit formula for this sequence is what you mean?

OpenStudy (anonymous):

yes @tHe_FiZiCx99

OpenStudy (the_fizicx99):

Is the sequence arithmetic or geometric?

OpenStudy (anonymous):

geometric I think! I got an = -3 + 4 (n-)

OpenStudy (anonymous):

@d3v13 no =/

OpenStudy (the_fizicx99):

One way to find out is by using the previous term, 9/5 = 1.8 5/1 = 5 and so on... They don't have a common ratio, where as if I subtracted 9-5 it gives me 4, 5-1 = 4 1-(-3) = 4 So they have a common ♦(difference)♦ so its arithmetic. The explicit formula for arithmetic is: an = a1 + d(n – 1) a1 = -3 d = 4 n = the term you want to find

OpenStudy (anonymous):

thank u so much for explaining!! :)))

OpenStudy (the_fizicx99):

yw